A distribution of values is normal with a mean of 197.3 and a standard deviation of 26.6. Find P9, which is the score separating the bottom 9% from the top 91%. P9 = 163 Incorrect Enter your answer as a number accurate to 1 decimal place. Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.
Answer #1

Solution :

mean = \mu = 197.3

standard deviation = \sigma = 26.6.

Using standard normal table,

P(Z > z) = 91%

1 - P(Z < z) = 0.09

P(Z < z) = 1 - 0.09 = 0.91

P(Z < 1.341) = 0.91

z =1.341

Using z-score formula,

x = z * \sigma + \mu

x = 1.341 * 26.6. + 197.3

   = 232.97

P9 = 233.0